C96L1

PI: Rioja/Sanghera
Title: CAL - DA193

Session: February 1996 (session 1)
Frequency: 18 cm
Mode: A
Obs time (UT): 40 EU 14:00(09/02)-14:30(09/02)
Obs time (UT): 45 EU 16:30(14/02)-17:30(14/02)
Stations: SLNRJWCaShUr
Processor: Bonn
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (foley at Fri Feb 16 05:49:43 EST 1996) Minor failures
Day 40 the first section (UT 1400-1413) the source is out of limits Day 45 OK Flux DA193 ~ 2.08Jy polarization <2%

Onsala (John Conway at Mon Feb 19 07:51:52 EST 1996) Minor failures
First part, phase cals may be large and increase system noise in a couple of channels Second Part. due to an error restarting schedule first two out of four tape passes were lost. Note that between first and second parts capstan was replaced as was capstan motor and head positions were recalibrated.

Simeiz (Alexandre Volvatch at Wed Feb 21 08:43:53 EST 1996) Success
On day 40 clock offset=0.6 microsec. Temp: 6 C; Press: 754 mm hg; Humid: 45 %; Cloudy. On day 45 clock offset=0.9 microsec Temp: 8 C; Press: 765 mm hg; Humid: 80-87%; Clear.

Urumqi (Lee Suqin at Wed Feb 21 09:15:17 EST 1996) Minor failures
On 9 Feb:cal temprature is 45k the system temperature is very high because there is one contactor in Lo being loosn. It is a fine day. Observation is OK. On 14 Feb: Cal temprature is 103k and the system temprature is about 210k. It is cloudy. Observation is OK.

Jodrell (Dave Henstock at Thu Feb 22 05:55:17 EST 1996) Minor failures
Bad RFI during both parts of experiment - probably worst in first part.. VCs 9,10 worst affected, but sporadic interference in VCs 4,7,11,12 too. Due to RFI, Tsys measurements only crudely calibrated. Specific Problems: Day 45 16:30 - 16:48 Field system crash - some or all of the data in this period was lost (some time on wrong pass). Tsys is in Jy. Need to multiply measured Tsys by 1.23 for better calibration.

Cambridge (Dave Henstock at Thu Feb 22 05:57:33 EST 1996) Success
No problems. Note that there are no Tsys measurements for Cambridge. Tsys is 212 Jy (from EVN status).

Medicina (Venturi Tiziana at Thu Feb 22 09:37:57 EST 1996) Success
Experiment carried out successfully. No problems reported.

Noto (Corrado Trigilio at Sat Feb 24 05:47:06 EST 1996) Minor failures
No phase cal for the 18cm receiver. Observations carried out successfully. Accidentally log file has been lost for first part of cal.

Shanghai (Huang Xinyong at Fri Mar 1 04:29:39 EST 1996) Success
All observations took place, but only in LCP. On day 40: clock offset = -1.0 usec, no phase cal signal. Weather was fine. On day 45: clock offset = -0.9 usec, with phase cal signal. Weather was overcast with strong win.